# The Limited Impact of Medical Adaptation of Large Language and Vision-Language Models

### v1

@@ 3 Experimental Setup
For visual QA, we use VQA-RAD (Lau et al., 2018), PathVQA (He et al., 2020), SLAKE (Liu et al., 2021), and MMMU-Medical (Yue et al., 2024) for evaluation. MMMU-Medical refers to a subset of MMMU corresponding to 5 subjects relevant to medicine: basic medical science, clinical medicine, diagnostics and laboratory medicine, pharmacy, and public health. For VQA-RAD, we address the train-test leakage and duplication issues in the official train–test splits, previously noted by Moor et al. (2023b), by removing the training examples repeated in the test set and removing all duplicates in both sets. We then take a random 80–20 split on the training set to create a new train–validation split, as the official split does not include a validation set. For MMMU-Medical, which does not have a public test set, we randomly select 5 examples from the official validation set for validation, and reserve the remaining 25 examples for testing. For all other datasets, we use the official split as provided. We provide the remaining details in Appendix A.
@@
Since we focus on closed-ended QA tasks, we use exact-match accuracy as our main evaluation metric. Following the Holistic Evaluation of Language Models (HELM) benchmark (Liang et al., 2023), we treat the text generated by a model (without any constraints on the vocabulary) to be its prediction, and check for an exact match between the prediction and the correct answer up to primitive string operations (e.g., lower-casing, removing white space/punctuation). To handle cases where the model simply repeats the list of answer choices or produces an ambiguous answer (e.g., selecting multiple answer choices), we take a conservative approach and treat the prediction to be incorrect, even if there is a match. Meanwhile, to quantify the extent of improvement from medical DAPT, we also consider the relative accuracy of the medical model with respect to the general-domain model. Formally, we define relative exact-match accuracy as $\mathbb{E}[\mathds{1}[f_{\text{medical}}(x)=y]-\mathds{1}[f_{\text{general}}(x)=y]]\in[-1,1]$, where $f_{\text{medical}}$ and $f_{\text{general}}$ denote the medical and general-domain models, $x$ and $y$ denote the input prompt and answer in a QA pair from the test set, and $\mathds{1}[\cdot]$ denotes the indicator function. This metric quantifies the difference in accuracy between the medical model and the general-domain model. To distinguish the two metrics, we refer to the former as the absolute exact-match accuracy in subsequent discussions.
@@ 3.1 Zero-/Few-shot Prompting with Model-Specific Prompt Selection
In this section, we provide an overview of our approach to assess whether medical DAPT leads to statistically significant improvements in zero-/few-shot medical QA performance. For few-shot prompting, we consider the 3-shot setting to ensure that the input prompt is shorter than the context window sizes for all models evaluated. Meanwhile, we exclude the EHRNoteQA and i2b2 datasets from few-shot prompting evaluations for medical LLMs, as most of the clinical notes in these datasets already occupy the full context window for most models even in the zero-shot setting. For evaluation, we pay special attention to two aspects. First, language models are highly sensitive to the choice of prompting strategy (e.g., prompt format, choice of few-shot examples), where seemingly insignificant changes to the prompt can lead to idiosyncratic model behavior (Jiang et al., 2020; Zhao et al., 2021; Ceballos-Arroyo et al., 2024b). Second, prior works show that the “optimal” choice of prompt format rarely correlates between different models (Sclar et al., 2024), suggesting that using a single, fixed prompt for all models for comparison can result in misleading conclusions.
Figure 2: Overview of the prompt format sampling (left) and model-specific prompt selection (right) process considered in our evaluations (see Section 3.1).
@@
To ensure a fair comparison that isolates the impact of medical DAPT, we treat the choice of prompt format and few-shot examples as additional hyperparameters when generating predictions, and tailor them to each model independently (Figure 2). We first randomly sample 10 plausible prompt formats from a predefined search space and 10 different sets of few-shot examples from the training set of each dataset. We then search over all pairs of prompt formats (plus one additional manually designed default format) and few-shot examples, and select the best pair out of $(10+1)\times 10=110$ that results in the highest validation exact-match accuracy. Given that a grid search at this scale can be computationally expensive, especially for datasets like MedMCQA that contain 37k validation QA pairs (see Table A1), we randomly subsample 500 validation QA pairs for datasets that have more than 500. Using the vLLM framework (Kwon et al., 2023) for sampling model outputs, this leads to a runtime of around 5–15 minutes per trial, when using 4 NVIDIA A6000 GPUs for the 70B models and 2 GPUs for the others. We then generate the final predictions on the test set using the prompt format and few-shot samples selected for each model. In the zero-shot setting, we only search over the prompt formats.
@@
To define the prompt format search space, we follow the approach by Sclar et al. (2024) and construct a context-free grammar of semantically equivalent yet syntactically distinct prompt formats (Figure 2, left). For the medical models that have a specific prompt format designed and recommended for closed-ended QA tasks (e.g., BioMistral (Labrak et al., 2024)), we fix the prompt format to what is provided and only search over the choice of few-shot examples. In the case when such information is missing or only partially available (see Table C1), we search over both the prompt formats and few-shot examples. For instruction-tuned models, which typically have a structured conversational format (e.g., ‘### User:…### Assistant:…”) that is expected, we use the sampled question and answer templates to format each “user” query and “assistant” response. We provide the remaining details in Appendices BC.
3.2 Supervised Fine-tuning
@@
To assess whether medical DAPT leads to a better initialization of the model parameters for supervised fine-tuning on downstream medical QA tasks, we compare the performances of models in each pair after fine-tuning each on the training split of each QA dataset. Due to computational constraints, we limit our discussions to LLM pairs with 7–8B parameters per model and all VLM pairs (which have 7–9B parameters per model) and focus on parameter-efficient fine-tuning instead of full fine-tuning. For textual QA evaluations, we exclude the MMLU-Medical datasets given that the training sets only include 5 examples per subject. For visual QA evaluations, we exclude the MMMU-Medical datasets for the same reason.
@@
For the Med-Flamingo-9B and Open-Flamingo-9B, we fully fine-tune the gated cross-attention layers, the Perceiver resampler (Jaegle et al., 2021), and the embeddings of the <image> and <|endofchunk|> special tokens [^4] See Section 3 of Awadalla et al. (2023) for more details on the architecture and special tokens., which already only correspond to a subset of all model parameters. For all other LLMs and VLMs, we fine-tune each model with low-rank adapters (LoRA; Hu et al., 2022). We add the low-rank adapters to all of the linear layers (i.e., the attention and the feedforward network layers in each Transformer block), following prior works that demonstrate its effectiveness over adapting a subset of the linear layers (Dettmers et al., 2023; Biderman et al., 2024). We fine-tune the trainable parameters by minimizing the cross-entropy loss on the output tokens generated conditional on the input context (i.e., the task instruction, image, and question). We train all models for a maximum of 10 epochs and apply early stopping regularization, monitoring the validation loss after every epoch of training and enforcing a patience of 1. To apply the gradient updates, we use the AdamW optimizer (Loshchilov and Hutter, 2019) with the recommended momentum hyperparameters $\beta_{1}=0.9,\beta_{2}=0.999$, and $\epsilon=10^{-8}$ (Kingma and Ba, 2015); and a cosine learning rate schedule with a linear warmup of 0.05 $\times$ maximum number of training steps. For Med-Flamingo-9B and Open-Flamingo-9B, we perform a grid search over the learning rates and weight decay coefficients and use the validation loss to select the best setting for final evaluation. For all other models, we perform a grid search over the LoRA ranks and learning rates and use the validation loss to select the best setting. We perform all experiments using the DeepSpeed Zero Redundancy Optimizer (ZeRO) (Rajbhandari et al., 2020) and PyTorch FSDP (Zhao et al., 2023) frameworks for distributed training. We provide the remaining details in Appendix D.
@@ 4 Zero-/Few-Shot Prompting Evaluation Results
Here, we summarize the main findings from the zero-/few-shot prompting experiments outlined in Section 3.1. Unless specified otherwise, we focus on the greedy decoding results in subsequent discussions and include the results for constrained decoding in Appendix E. Overall, we find that all medical VLMs and the majority of medical LLMs fail to consistently improve over their general-domain counterparts in the zero-/few-shot prompting regime (Section 4.1). Notably, we find that the medical LLMs do not show consistent improvements on either the medical knowledge QA datasets or the clinical note QA datasets in this setting. Moreover, we demonstrate the importance of rigorous experimental design in surfacing this finding—performing pairwise model comparison with a single, fixed prompt optimized only for the medical model, while ignoring statistical uncertainty, can paint a misleadingly optimistic picture of off-the-shelf performance benefits from medical DAPT (Section 4.2).
@@
We first show that all medical VLMs and the majority of medical LLMs fail to consistently outperform their corresponding base models in terms of zero-shot and 3-shot performance, after model-specific prompt selection and statistical testing. We calculate the 95% confidence intervals in relative exact-match accuracy via bootstrapping on the test set, as described in Section 3. Below, we present the results on textual medical knowledge QA datasets, textual clinical note QA datasets, and the visual QA datasets one-by-one.
Figure 3: After independently selecting the best prompt format and examples for each model, medical LLMs do not consistently show a statistically significant improvement over their general-domain counterparts on textual medical knowledge QA tasks, in both (a) zero-shot and (b) 3-shot settings. The top row of each panel shows the absolute exact-match accuracies on the test set, and the bottom row of each panel shows the relative exact-match accuracies along with 95% confidence intervals derived via bootstrapping on the test set (see Section 3). Here, we show the results when model predictions are generated via greedy decoding. The improvements are also limited with constrained decoding (see Figure E1).
@@ 4.1.1 Evaluation of Medical LLMs on Textual Medical Knowledge QA
In Figure 3, we show the absolute and relative exact-match accuracies achieved by the medical and general-domain LLMs on the textual medical knowledge QA datasets, from zero-shot (Figure 3(a)) and 3-shot prompting (Figure 3(b)), respectively. In Table 2, we also show the win, tie, and loss rates (%) of the medical LLMs, where win rate refers to the proportion of QA datasets where a medical model shows a statistically significant improvement over its base model. For each medical model in Table 2, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. We exclude the results for Clinical-Camel-70B on both versions of MedQA, and the Med42 models on all textual medical knowledge QA datasets, as they have already been trained on these datasets as part of medical DAPT (see Table 1).
The top rows of Figure 3(a) and (b) show that the absolute exact-match accuracies are mostly similar between each model pair across all datasets and model scales, with marginal performance improvements. Going from the zero-shot setting to the 3-shot setting, we observe that both medical and general-domain models generally achieve higher accuracy across all datasets, but the pairwise differences in performance remain similarly small.
We also find that more recent general-domain models tend to show stronger performance overall even without any medical adaptation via DAPT. For example, Llama-3-8B (orange) performs on par with the medically adapted MediTron-70B and Clinical-Camel-70B (red) in both zero-shot and 3-shot settings, despite the substantial difference in model scale.
The bottom rows of Figures 3(a) and (b) more clearly show that in terms of relative exact-match accuracy, the majority of medical models show little to no improvement over their general-domain base models, with the 95% confidence intervals crossing or lying below zero. In fact, the win, tie, and loss rates shown in Table 2 show that only 2 out of 7 medical LLMs show statistically significant improvements in performance (i.e., win more than they lose to their base models) in both zero-shot and 3-shot settings. In the zero-shot setting, only Clinical-Camel-70B (red) and BioMistral-7B (green) show improvements, each achieving win/tie/loss rates of 27.3%/63.6%/9.1% and 30.8%/69.2%/0%, respectively. In the 3-shot setting, only OpenBioLLM-70B (blue) and BioMistral-7B (green) show improvements, each achieving win/tie/loss rates of 30.8%/69.2%/0% and 46.2%/53.8%/0%, respectively. Notably, we find that BioMistral-7B is the only medical LLM that consistently outperforms its base model (Mistral-7B-Instruct-v0.1) across both zero-shot and 3-shot settings, albeit with relatively low absolute performance.
Meanwhile, we observe that some medical LLMs (e.g., MediTron-7B and BioMedGPT-LM-7B) actually perform significantly worse than their base models, with loss rates of 76.9% and 92.3% in the 3-shot setting. When we aggregate the results over all (model pair, textual medical knowledge QA dataset) combinations (last row of Table 2), we find that the medical LLMs achieve win/tie/loss rates of 9.4%/56.3%/34.3% in the zero-shot setting and 12.1%/49.8%/38.2% in the 3-shot setting, indicating that the medical LLMs overall reach a statistical tie with their base models in most cases and only win in a small number of cases.
We similarly observe limited improvements overall with constrained decoding (see Appendix E). As shown in Figure E1 and Table E1, when we aggregate the results over all (model pair, textual medical knowledge QA dataset) combinations, medical LLMs achieve win/tie/loss rates of 16.9%/68.6%/14.5% in the zero-shot setting and 11.2%/74.1%/14.7% in the 3-shot setting. While the aggregate loss rates are generally lower with constrained decoding, the majority of cases still result in a tie, and the win rates are roughly equivalent to the loss rates. Meanwhile, we observe that some medical LLMs show larger improvements with constrained decoding. For example, when switching from greedy decoding to constrained decoding, we observe that the zero-shot win/tie/loss rates for MediTron-70B change from 0%/61.5%/38.5% to 30.8%/46.2%/23.1%, showing a substantial increase in the win rate. However, the results are mixed, as some other models (e.g., Clinical-Camel-70B and OpenBioLLM-70B) perform worse with constrained decoding.
In summary, these results show that after appropriately accounting for prompt sensitivity and statistical uncertainty, the medical LLMs trained via DAPT show limited improvements in zero-/few-shot prompting performance over their general-domain counterparts, on textual QA tasks focused on evaluating medical knowledge.
@@
Table 2: The zero-shot and 3-shot win/tie/loss rates (%) for all medical LLMs on textual medical knowledge QA, after independently optimizing the prompt for each model. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. Here, we show the results when model predictions are generated via greedy decoding. The results for constrained decoding are shown in Table E1.
@@
OpenBioLLM-70B (Pal and Sankarasubbu, 2024) 7.7 69.2 23.1 30.8 69.2 0
MediTron-70B (Chen et al., 2023) 0 61.5 38.5 0 69.2 30.8
Clinical-Camel-70B (Toma et al., 2023) 27.3 63.6 9.1 0 63.6 36.4
OpenBioLLM-8B (Pal and Sankarasubbu, 2024) 0 46.2 53.8 7.7 61.5 30.8
MediTron-7B (Chen et al., 2023) 0 69.2 30.8 0 23.1 76.9
BioMistral-7B (Labrak et al., 2024) 30.8 69.2 0 46.2 53.8 0
BioMedGPT-LM-7B (Luo et al., 2023) 0 15.4 84.6 0 7.7 92.3
Aggregate 9.4 56.3 34.3 12.1 49.8 38.2
Figure 4: After independently selecting the best prompt format and examples for each model, medical LLMs do not consistently show a statistically significant improvement over their general-domain counterparts on textual clinical note QA tasks, in both (a) zero-shot and (b) 3-shot settings. In the 3-shot setting, we exclude the results on the EHRNoteQA and i2b2 datasets due to the limited context window sizes of the medical LLMs that we evaluate (see Section 3.1). The top row of each panel shows the absolute exact-match accuracies on the test set, and the bottom row of each panel shows the relative exact-match accuracies along with 95% confidence intervals derived via bootstrapping on the test set (see Section 3). Here, we show the results when model predictions are generated via greedy decoding. The improvements are also limited with constrained decoding (see Figure E2).

### v2

@@ 3 Experimental Setup
For visual medical QA, we use VQA-RAD (Lau et al., 2018), PathVQA (He et al., 2020), SLAKE (Liu et al., 2021), and MMMU-Medical (Yue et al., 2024) for evaluation. MMMU-Medical is a subset of MMMU with 5 subjects relevant to medicine: basic medical science, clinical medicine, diagnostics and lab medicine, pharmacy, and public health. For VQA-RAD, we address the train–test leakage and duplication issues in the official train–test split (Moor et al., 2023b) by removing the training examples repeated in the test set and removing all duplicates in both sets. As the official split does not include a validation set, we take a random 80–20 split on the training set to create a new train–validation split. For MMMU-Medical, which lacks a public test set, we randomly select 5 examples from the official validation set for validation, and reserve the remaining 25 examples for testing. For all other datasets, we use the official split as provided.
@@
Since we focus on closed-ended QA tasks, we use exact-match accuracy as our main evaluation metric. Following the Holistic Evaluation of Language Models (HELM) benchmark (Liang et al., 2023), we treat the text generated by a model (without any constraints on the vocabulary) to be its prediction, and check for an exact match between the prediction and the correct answer up to primitive string operations (e.g., lower-casing, removing white space/punctuation). If multiple matches occur, we treat the prediction as incorrect (even if one of the matches is the correct answer), in order to handle cases where the model simply repeats the list of answer choices or produces an ambiguous answer (e.g., selecting multiple answer choices). Meanwhile, to quantify the extent of improvement from medical DAPT, we also consider the relative accuracy of the medical model with respect to the general-domain model. Formally, we define relative exact-match accuracy as $\mathbb{E}[\mathds{1}[f_{\text{medical}}(x)=y]-\mathds{1}[f_{\text{general}}(x)=y]]\in[-1,1]$, where $f_{\text{medical}}$ and $f_{\text{general}}$ denote the medical and general-domain models, $x$ and $y$ denote the input prompt and answer in a QA pair from the test set, and $\mathds{1}[\cdot]$ denotes the indicator function. This metric quantifies the difference in accuracy between the medical model and the general-domain model. We refer to the accuracy of a particular model as the absolute exact-match accuracy in subsequent discussions, and refer to the metric above (for a pair of models) as the relative accuracy.
@@ 3.1 Zero-/Few-shot Prompting with Model-Specific Prompt Selection
In this section, we provide an overview of our approach to assess whether medical DAPT leads to statistically significant improvements in zero-/few-shot medical QA performance. For few-shot prompting, we consider the 3-shot setting to ensure that the input prompt is shorter than the context window sizes for all models evaluated. Meanwhile, we exclude the EHRNoteQA and i2b2 datasets from few-shot prompting evaluations for medical LLMs, as most of the clinical notes in these datasets already occupy the full context window for most models even in the zero-shot setting. For evaluation, we pay special attention to two aspects. First, language models are highly sensitive to the choice of prompting strategy (e.g., prompt format, choice of few-shot examples), where seemingly insignificant changes to the prompt can lead to idiosyncratic model behavior (Jiang et al., 2020; Zhao et al., 2021; Ceballos-Arroyo et al., 2024). Second, prior works show that the “optimal” choice of prompt format is rarely the same between different models (Sclar et al., 2024), suggesting that using a single, fixed prompt for all models for comparison can result in misleading conclusions.
@@
To ensure a fair comparison that isolates the impact of medical DAPT, we treat the choice of prompt format and few-shot examples as additional hyperparameters when generating predictions, and tailor them to each model independently (Figure 3). We first randomly sample 10 plausible prompt formats from a predefined search space and 10 different sets of few-shot examples from the training set of each dataset. We then search over all pairs of prompt formats (plus one additional manually designed default format) and few-shot examples, and select the best pair out of $(10+1)\times 10=110$ that results in the highest validation exact-match accuracy. Given that a grid search at this scale can be computationally expensive, especially for datasets like MedMCQA that contain 37k validation QA pairs (Table A1), we randomly subsample 500 validation QA pairs for datasets that have more than 500. Using the vLLM framework (Kwon et al., 2023) for sampling model outputs, this leads to a runtime of around 5–15 minutes per trial, when using 4 NVIDIA A6000 GPUs for the 70B models and 2 GPUs for the others. We then generate the final predictions on the test set using the prompt format and few-shot samples selected for each model. In the zero-shot setting, we only search over the prompt formats. Meanwhile, we acknowledge that in the “truly” zero-shot setting, it may not be possible to optimize the prompt with respect to a dedicated validation set. Nonetheless, we implement our prompt selection approach even in the zero-shot setting to mitigate the potential impact that choosing a particular prompt format has on QA performance, and ensure a fair comparison between models in each pair.
@@
To define the prompt format search space, we follow Sclar et al. (2024) and construct a context-free grammar of semantically equivalent yet syntactically distinct prompt formats (Figure 3, left). For the medical models that have a specific prompt format designed and recommended for closed-ended QA tasks (e.g., BioMistral (Labrak et al., 2024)), we fix the prompt format to what is provided and only search over the choice of few-shot examples. When such information is missing or only partially available (Table C1), we search over both the prompt formats and few-shot examples. For instruction-tuned models, which typically have a structured conversational format that is expected (e.g., “### User:…### Assistant:…”), we use the sampled question and answer templates to format each “user” query and “assistant” response. We provide the remaining details in Appendices BC.
Figure 3: Overview of the prompt format sampling (left) and model-specific prompt selection (right) process considered in our evaluations (Section 3.1).
3.2 Supervised Fine-Tuning (SFT)
@@
To assess whether medical DAPT leads to a better initialization of the model parameters for fine-tuning on downstream medical QA tasks, we compare the performances of models in each pair after fine-tuning each on the training set of each QA dataset. Due to computational constraints, we focus on parameter-efficient fine-tuning methods (detailed below) instead of full fine-tuning. For the textual medical QA evaluations, we exclude the MMLU-Medical datasets given that the training sets only include 5 examples per subject. For the visual medical QA evaluations, we exclude the MMMU-Medical datasets for the same reason.
@@
For Med-Flamingo-9B and Open-Flamingo-9B, we fine-tune the gated cross-attention layers, the Perceiver resampler (Jaegle et al., 2021), and the embeddings of the special <image> and <|endofchunk|> tokens, which correspond to a only subset of all model parameters. For all other LLMs and VLMs, we fine-tune each model with low-rank adapters (LoRA; Hu et al., 2022). We add the low-rank adapters to all of the linear layers (i.e., the self-attention and feedforward network layers in each Transformer block), following prior works that demonstrate its effectiveness over adapting a subset of the linear layers (Dettmers et al., 2023; Biderman et al., 2024). We fine-tune the trainable parameters by minimizing the cross-entropy loss on the output tokens generated conditional on the input context (i.e., the task instruction, image, and question). We train all models for a maximum of 10 epochs and apply early stopping regularization, monitoring the validation loss after every epoch of training and enforcing a patience of 1. To apply the gradient updates, we use the AdamW optimizer (Loshchilov and Hutter, 2019) with the recommended momentum hyperparameters $\beta_{1}=0.9,\beta_{2}=0.999$, and $\epsilon=10^{-8}$ (Kingma and Ba, 2015); and a cosine learning rate schedule with a linear warmup of 0.05 $\times$ maximum number of training steps. For Med-Flamingo-9B and Open-Flamingo-9B, we perform a grid search over the learning rates and weight decay coefficients and use the validation loss to select the best setting for final evaluation. For all other models, we perform a grid search over the LoRA ranks and learning rates and use the validation loss to select the best setting. We perform all experiments using the DeepSpeed ZeRO (Rajbhandari et al., 2020) and PyTorch FSDP (Zhao et al., 2023) frameworks for distributed training. Meanwhile, for fine-tuning the 70B-parameter LLMs on PubMedQA and i2b2 (Diabetes), we also apply 4-bit NormalFloat double quantization (QLoRA; Dettmers et al., 2023) to the model parameters to address out-of-memory issues. We provide the remaining details in Appendix E.
@@ 4 Zero-/Few-Shot Prompting Evaluation Results
We summarize the main findings from the zero-/few-shot prompting experiments outlined in Section 3.1. Unless specified otherwise, we focus on the greedy decoding results in subsequent discussions and include the results for constrained decoding in Appendix D. Overall, we find that all medical VLMs and the majority of medical LLMs fail to consistently improve over their general-domain counterparts in the zero-/few-shot prompting regime (Section 4.1). We find that the medical LLMs do not show consistent improvements on either medical knowledge QA or clinical note QA datasets in this setting. Moreover, we demonstrate the importance of rigorous experimental design in surfacing this finding—performing pairwise model comparison with a single, fixed prompt optimized only for the medical model, while ignoring statistical uncertainty, can paint a misleadingly optimistic picture of off-the-shelf performance benefits from medical DAPT (Section 4.2).
@@
We first show that all medical VLMs and the majority of medical LLMs fail to consistently outperform their corresponding base models in terms of zero-shot and 3-shot performance, after model-specific prompt selection and statistical testing. We calculate the 95% confidence intervals in relative exact-match accuracy via bootstrapping on the test set, as described in Section 3. Below, we present the results on the textual medical knowledge QA datasets, the textual clinical note QA datasets, and the visual medical QA datasets one-by-one.
Figure 4: After independently selecting the best prompt format and examples for each model, medical LLMs (textured bars) fail to consistently improve over their base models (solid bars) on textual medical knowledge QA tasks, in both (a) zero-shot and (b) 3-shot settings. In each panel, the top row shows the absolute exact-match accuracies on the test set, and the bottom row shows the relative exact-match accuracies along with 95% confidence intervals derived via bootstrapping on the test set (Section 3). Here, model predictions are generated via greedy decoding. Improvements are also limited with constrained decoding (Figure D1).
@@ 4.1.1 Evaluation of Medical LLMs on Textual Medical Knowledge QA
In Figures 4(a) and (b), we show the absolute and relative exact-match accuracies achieved by the medical and general-domain LLMs on the textual medical knowledge QA datasets, from zero-shot and 3-shot prompting, respectively. In Table 3, we show the win, tie, and loss rates (%) of the medical LLMs, where win rate refers to the proportion of QA datasets where a medical model shows a statistically significant improvement over its base model. We exclude the results for Clinical-Camel-70B on both versions of MedQA, and the Med42 models on all textual medical knowledge QA datasets, as they have already been trained on these datasets as part of medical DAPT (Table 2).
Overall, we find that the majority of medical models show little to no improvement over their base models, with the 95% confidence intervals in relative accuracy crossing or lying below zero (bottom rows of Figure 4(a) and (b)).
@@
Table 3: The zero-shot and 3-shot win/tie/loss rates (%) for all medical LLMs on textual medical knowledge QA, after independently optimizing the prompt for each model. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. Here, we show the results when model predictions are generated via greedy decoding. The results for constrained decoding are shown in Table D1.
@@
OpenBioLLM-70B 7.7 69.2 23.1 30.8 69.2 0
MediTron-70B 0 61.5 38.5 0 69.2 30.8
Clinical-Camel-70B 27.3 63.6 9.1 0 63.6 36.4
OpenBioLLM-8B 0 46.2 53.8 7.7 61.5 30.8
MediTron-7B 0 69.2 30.8 0 23.1 76.9
BioMistral-7B 30.8 69.2 0 46.2 53.8 0
BioMedGPT-LM-7B 0 15.4 84.6 0 7.7 92.3
Aggregate 9.0 56.2 34.8 12.4 49.4 38.2
In fact, in each setting (zero-shot and 3-shot), only 2 out of 7 medical LLMs show statistically significant improvements in performance (i.e., win more than they lose to their base models), and only one medical LLM shows statistically significant improvements in both settings (Table 3). In the zero-shot setting, only Clinical-Camel-70B (orange) and BioMistral-7B (purple) show improvements, each achieving win/tie/loss rates of 27.3%/63.6%/9.1% and 30.8%/69.2%/0%, respectively. In the 3-shot setting, only OpenBioLLM-70B (blue) and BioMistral-7B (purple) show significant improvements, each achieving win/tie/loss rates of 30.8%/69.2%/0% and 46.2%/53.8%/0%, respectively. Notably, BioMistral-7B is the only medical LLM that consistently outperforms its base model (Mistral-7B-Instruct-v0.1) across both settings, albeit with relatively low absolute performance.
Meanwhile, we observe that some medical LLMs actually perform significantly worse than their base models—e.g., MediTron-7B and BioMedGPT-LM-7B with loss rates of 76.9% and 92.3% in the 3-shot setting. When we aggregate the results over all (model pair, textual medical knowledge QA dataset) combinations (last row of Table 3), we find that the medical LLMs achieve win/tie/loss rates of 9.0%/56.2%/34.8% in the zero-shot setting and 12.4%/49.4%/38.2% in the 3-shot setting, indicating that the medical LLMs overall reach a statistical tie with their base models in most cases and only win in a small number of cases.
We similarly observe limited improvements overall with constrained decoding (Appendix D.1). When we aggregate the results over all (model pair, textual medical knowledge QA dataset) combinations, medical LLMs achieve win/tie/loss rates of 16.9%/68.5%/14.6% in the zero-shot setting and 11.2%/74.2%/14.6% in the 3-shot setting (Table D1). While the aggregate loss rates are generally lower with constrained decoding, the majority of cases still result in a tie, with similar aggregate win and loss rates. Meanwhile, we observe that some medical LLMs show larger improvements with constrained decoding. For example, when switching from greedy to constrained decoding, the zero-shot win/tie/loss rates for MediTron-70B change from 0%/61.5%/38.5% to 30.8%/46.2%/23.1%, showing a substantial increase in the win rate. However, the results are mixed, as some other models (e.g., Clinical-Camel-70B, OpenBioLLM-70B) perform worse with constrained decoding.
In summary, these results show that after appropriately accounting for prompt sensitivity and statistical uncertainty, medical LLMs trained via DAPT show limited improvements in zero-/few-shot prompting performance over their general-domain base models, on textual QA tasks focused on evaluating medical knowledge.